# A Bayesian approach to functional regression: theory and computation

### v2

@@ 3 Posterior consistency
This section demonstrates how the RKHS-based Bayesian methodology proposed herein establishes a robust theoretical foundation for subsequent prediction procedures. Firstly, let us briefly recall what we understand by posterior consistency. Note that to avoid confusion, throughout this section we will sometimes use bold letters to represent random variables. Consider a sample space $\mathcal{X}$ and $X_{1},\dots,X_{n}$ an i.i.d. sample of the data $X$. Let us fix a prior distribution $\Pi$ for random variables $\bm{\theta}$ on the parameter space $\Theta$, that is, $\bm{\theta}\sim\Pi$, and let $P_{\theta}$ represent a sampling model (a distribution on $\mathcal{X}$ indexed by $\theta\in\Theta$) such that $X|\bm{\theta}\sim P_{\bm{\theta}}$. Furthermore, assume that the model is well-specified, i.e., there is a true value $\theta_{0}\in\Theta$ such that $X\sim P_{\theta_{0}}$, and denote by $P_{0}^{\infty}$ the joint probability measure of $(X_{1},X_{2},\dots)$ when $\theta_{0}$ is the true value of the parameter.
Definition 1 (Ghosh and Ramamoorthi, 2003).
@@
We say that the posterior distribution is (strongly) consistent at $\theta_{0}$ if for every neighborhood $U$ of $\theta_{0}$ it holds that
[…]
For a metric space $(\Theta,d)$, this is equivalent to
$\lim_{n\to\infty}\Pi(d(\bm{\theta},\theta_{0})<\varepsilon\mid X_{1},\dots,X_{n})=1\quad P_{0}^{\infty}-\text{a.s.},\quad\text{for all }\varepsilon>0.$
@@
Note that the conditional probabilities are computed under the assumed joint distribution of $(\bm{\theta},(X_{1},X_{2},\dots))$. Essentially, we are saying that the posterior concentrates around $\theta_{0}$ for almost all sequences of data. Thus, if consistency holds, the effect of the prior gets diluted as more and more data is used for the inference.
Doob’s theorem
It turns out that, under very general conditions, the posterior distribution is consistent at almost every value of $\theta_{0}$ with respect to the measure induced by the prior (Doob, 1949). Let the sample space $\mathcal{X}$ and the parameter space $\Theta$ be complete separable metric spaces, endowed with their respective Borel sigma-algebras. Consider the model $\bm{\theta}\sim\Pi$ and $X_{1},X_{2},\ldots|\bm{\theta}\sim P_{\bm{\theta}}$ i.i.d., where $\bm{\theta}$ is a random variable taking values in $\Theta$. Observe that this induces a posterior distribution $\Pi(\bm{\theta}|X_{1},\dots,X_{n})$.
@@
If $\theta\mapsto P_{\theta}$ is one-to-one and $\theta\mapsto P_{\theta}(A)$ is measurable for all measurable sets $A\subseteq\mathcal{X}$, then the posterior distribution is consistent at $\Pi$-almost all values of $\Theta$. That is, there exists $\Theta_{*}\subseteq\Theta$ such that $\Pi(\Theta_{*})=1$ and for all $\theta_{0}\in\Theta_{*}$, if $X_{1},X_{2},\ldots\sim P_{\theta_{0}}$ i.i.d., then for any neighborhood $B$ of $\theta_{0}$ we have
[…]
See chapter 1.3 of Ghosh and Ramamoorthi (2003) for more details on this result. As a side note, in a nonparametric setting where the object of interest is a random function (e.g. a probability density), there is also a stronger consistency result by Schwartz (1965) which omits the $\Pi$-almost sure qualification under some more restrictive conditions. Moreover, there are some extensions of this result that deal with independent but not identically distributed data (Choi and Ramamoorthi, 2008).
3.1 Consistency in our RKHS model
Now we study in detail how Doob’s theorem can be applied in our linear RKHS model, but the posterior consistency results we obtain hold mutatis mutandis in the logistic case. In this section we will assume that the covariance function $K$ of the underlying stochastic process $X$ is strictly positive definite, which is not a very restrictive condition in practice. In the linear case the sample space is $\mathcal{X}\times\mathcal{Y}=L^{2}[0,1]\times\mathbb{R}$, which is already a complete separable metric space. Next, consider for each $p\in\mathbb{N}$ the subset of the $(2p+2)$-Euclidean space $\Theta_{p}=\{(b,\tau,\alpha_{0},\sigma^{2}):b=(\beta_{1},\dots,\beta_{p})\in\mathbb{R}^{p},\tau=(t_{1},\dots,t_{p})\in[0,1]^{p},\alpha_{0}\in\mathbb{R},\sigma^{2}\in\mathbb{R}^{+}_{0}\}$. Then, we can write our infinite-dimensional parameter space as
$\Theta=\bigcup_{p=1}^{\infty}\Theta_{p}.$ (3.1)
@@ 3.1 Consistency in our RKHS model
At this point we can follow an approach very similar to the one in Miller (2023), where posterior consistency is established in a mixture model with an unknown number of components, which has an infinite-dimensional parameter space that factorizes in the same way as (3.1). Note that given $\theta\in\Theta$ there is a unique $p=p(\theta)$ such that $\theta\in\Theta_{p}$. Considering that we will only be interested in small balls around the true value of the parameter, we can define a metric for $\theta,\theta^{\prime}\in\Theta$ by
@@
$d(\theta,\theta^{\prime})=\begin{cases}\min\left\{\|\theta-\theta^{\prime}\|,1\right\},\quad&\text{if }p(\theta)=p(\theta^{\prime}),\\ 1,\quad&\text{otherwise}.\end{cases}$ (3.2)
Since each $\Theta_{p}$ is itself a complete separable metric space with the inherited Euclidean norm, Proposition A.1 in Miller (2023) ensures that $(\Theta,d)$ is a complete separable metric space. Further, we equip both $\mathcal{X}\times\mathcal{Y}$ and $\Theta$ with their respective Borel sigma-algebras.
In terms of $\theta\in\Theta$, the data distribution is $P_{\theta}(X,Y)=P_{b,\tau,\alpha_{0},\sigma^{2}}(X,Y)$, and formally the joint distribution $P_{\theta}(X,Y)$ factorizes as $P_{\theta}(X,Y)=P(X)P_{\theta}(Y|X)$. Here $P(X)$ is the distribution of the underlying process $X$, and in our RKHS setting, $P_{\theta}(Y|X)\equiv\mathcal{N}(\alpha_{0}+\sum_{j=1}^{p(\theta)}\beta_{j}X(t_{j}),\,\sigma^{2})$. Now, let us suppose that $\theta\mapsto P_{\theta}(X,Y)(A)$ is measurable for all measurable sets $A\subseteq\mathcal{X}\times\mathcal{Y}$, which is true under mild conditions (see Appendix B). Moreover, for convenience, we will denote the sequences $(X,Y)_{1:n}=(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ and $(X,Y)_{1:\infty}=(X_{1},Y_{1}),(X_{2},Y_{2}),\dots$.
Now, the full hierarchical model under consideration is
(no. of components) $\mathcal{P}\sim\pi,$ (3.3)
(component values) $\bm{b}\mid\mathcal{P}=p\sim F_{p},$
(component times) $\bm{\tau}\mid\mathcal{P}=p\sim G_{p},$
(intercept) $\bm{\alpha_{0}}\sim C,$
(error variance) $\bm{\sigma^{2}}\sim D,$
[…]
@@
where $\pi$, $F_{p}$, $G_{p}$, $C$ and $D$ are probability measures on $\mathbb{N}$, $\mathbb{R}^{p}$, $[0,1]^{p}$, $\mathbb{R}$ and $\mathbb{R}^{+}_{0}$, respectively. Note that since $P_{\theta}(X,Y)$ is invariant under permutations of the component labels $b$ and $\tau$, we can only show consistency up to one such permutation. To that effect, and mirroring the strategy in Miller (2023), let $S_{p}$ denote the set of permutations of $\{1,\dots,p\}$, and for $\nu\in S_{p}$ and $\theta\in\Theta_{p}$, denote by $\theta[\nu]$ the result of applying the permutation $\nu$ to the component labels of $\theta$. That is, if $\theta=(\beta_{1},\dots,\beta_{p},t_{1},\dots,t_{p},\alpha_{0},\sigma^{2})$, then $\theta[\nu]=(\beta_{\nu_{1}},\dots,\beta_{\nu_{p}},t_{\nu_{1}},\dots,t_{\nu_{p}},\alpha_{0},\sigma^{2})$. Next, define $\tilde{B}(\theta_{0},\varepsilon)=\bigcup_{\nu\in S_{p}}\{\theta\in\Theta:d(\theta,\theta_{0}[\nu])<\varepsilon\}$ for $\theta_{0}\in\Theta_{p}$ and $\varepsilon>0$, which is the set of all parameters that are within $\varepsilon$ of some permutation of (the component labels of) $\theta_{0}$. Lastly, define the random variable $\bm{\theta}=(\bm{b,\tau,\alpha_{0},\sigma^{2}})$, which takes values in $\Theta$, and denote by $\Pi$ the prior distribution on $\bm{\theta}$ implied by the model in (3.3). In order for identifiability to hold, we need to place some restrictions on the prior:
Condition 3 (Identifiability constraints).
@@
Both assumptions can be interpreted as a way of pursuing parsimony in the model, aiming for as few components as possible. In practical and computational terms, we can think of $\delta$ as the machine precision number, so that virtually all continuous prior distributions satisfy the associated condition when implemented numerically in a computer. With this setup in mind, we are now ready to state our main consistency result:
@@
Suppose that Condition 3 holds. Then there exists $\Theta_{*}\subseteq\Theta$ such that $\Pi(\bm{\theta}\in\Theta_{*})=1$ and for all $\theta_{0}\in\Theta_{*}$, if $(X,Y)_{1:\infty}\sim P_{\theta_{0}}(X,Y)$ i.i.d., then for all $\varepsilon>0$
[…]
@@
[…]

### v3

@@ 3 Posterior consistency
This section explores the theoretical foundations of the proposed Bayesian models in the context of predictive inference. In particular, we establish results grounded in the theory of posterior consistency, which provides rigorous justification for the reliability of these models in asymptotic settings. For an in-depth treatment of posterior consistency and related concepts, we refer the reader to Ghosal and van der Vaart (2017).
@@
Firstly, let us recall what we understand by posterior consistency and posterior contraction rates. Consider the general setting of an i.i.d. sample $X_{1},\dots,X_{n}$ from a random variable $X$ taking values in a certain space $\mathcal{X}$. Let us fix a prior distribution $\Pi$ for random variables $\theta$ on the parameter space $\Theta$, that is, $\theta\sim\Pi$, and let $P_{\theta}$ represent a sampling model (a distribution on $\mathcal{X}$ indexed by $\theta\in\Theta$) such that $X|\theta\sim P_{\theta}$. Furthermore, assume that the model is well-specified, i.e., there is a true value $\theta_{0}\in\Theta$ such that $X\sim P_{\theta_{0}}$, and denote by $P_{0}^{\infty}$ and $P_{0}^{(n)}$ the joint probability measure of $(X_{1},X_{2},\dots)$ and $(X_{1},X_{2},\dots,X_{n})$, respectively, when $\theta_{0}$ is the true value of the parameter.
Definition 1.
@@
We say that the posterior distribution is (strongly) consistent at $\theta_{0}$ if for every neighborhood $B$ of $\theta_{0}$ (which for a metric space can just be the open balls around $\theta_{0}$) we have
$\lim_{n\to\infty}\Pi_{n}(\theta\in B\mid X_{1},\dots,X_{n})=1\quad P_{0}^{\infty}-\text{a.s.}$
@@
Note that the conditional probabilities are computed under the assumed joint distribution of $(\theta,(X_{1},X_{2},\dots))$. Essentially, we are saying that the posterior concentrates around $\theta_{0}$ for almost all sequences of data, and thus the effect of the prior gets diluted as more and more data is available for the inference. Furthermore, when the parameter space carries a metric, say $d$, we are also interested in the speed at which the posterior approaches the true parameter $\theta_{0}$. This is what we call a posterior contraction rate, which is a significant refinement of the comparatively weaker concept of consistency.
Definition 2.
A sequence $\varepsilon_{n}\to 0$ is a posterior contraction rate at $\theta_{0}$ with respect to $d$ if for every $M_{n}\to\infty$ the posterior satisfies $\Pi_{n}(\theta:d(\theta_{0},\theta)\geq M_{n}\varepsilon_{n}|X_{1},\dots,X_{n})\to 0$ in $P_{0}^{(n)}$-probability.
Naturally, once we have established a contraction rate, every slower sequence is also a valid contraction rate. Even though we are interested in the fastest possible rate, this is often difficult to compute or does not exist at all. Hence, we are usually content with finding a good enough rate, which we call “the” rate of contraction for the model.
In the rest of the section we analyze how these frequentist concepts apply to our Bayesian functional models; we focus on the linear case, but the results we obtain generally hold mutatis mutandis in the logistic case. All technical details and proofs are deferred to Appendix B. For the functional linear model, our data is an i.i.d. sample $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of a random vector $(X,Y)$, where $X=X(t)$ is a second-order stochastic process taking values in $\mathcal{X}=\mathcal{C}[0,1]$, the space of continuous functions, and $Y$ is a random variable taking values in $\mathcal{Y}=\mathbb{R}$. The state space is $\mathcal{X}\times\mathcal{Y}=\mathcal{C}[0,1]\times\mathbb{R}$, which is a complete separable metric space. Meanwhile, the infinite-dimensional parameter space can be characterized via the Euclidean spaces $\Theta_{p}=\{(b,\tau,\alpha,\sigma^{2}):b=(\beta_{1},\dots,\beta_{p})\in\mathbb{R}^{p},\tau=(t_{1},\dots,t_{p})\in[0,1]^{p},\alpha\in\mathbb{R},\sigma^{2}\in\mathbb{R}^{+}_{0}\}$ as the infinite union
$\Theta=\bigcup_{p=1}^{\infty}\Theta_{p},$ (3.1)
and note that given $\theta\in\Theta$ there is a unique $p=p(\theta)$ such that $\theta\in\Theta_{p}$. As usual, we equip both $\mathcal{X}\times\mathcal{Y}$ and $\Theta$ with their respective Borel sigma-algebras.
In terms of $\theta\in\Theta$, the data distribution can be written as $P_{\theta}(X,Y)=P_{b,\tau,\alpha,\sigma^{2}}(X,Y)$, and formally the joint distribution factorizes as $P_{\theta}(X,Y)=Q_{X}P_{\theta}(Y|X)$. Here $Q_{X}$ is the distribution of the underlying process $X$, and in our RKHS setting, $P_{\theta}(Y|X)$ is the normal distribution $\mathcal{N}(\alpha+\sum_{j=1}^{p(\theta)}\beta_{j}X(t_{j}),\,\sigma^{2})$. Moreover, for convenience, we will denote the sequences $(X,Y)_{1:\infty}:=(X_{1},Y_{1}),(X_{2},Y_{2}),\dots$ and $(X,Y)_{1:n}:=(X_{1},Y_{1}),(X_{2},Y_{2}),\dots,(X_{n},Y_{n})$. The full hierarchical model under consideration is
(no. of components) $\mathcal{P}\sim\pi,$ (3.2)
(component values) $b\mid\mathcal{P}=p\sim\phi_{p},$
(component times) $\tau\mid\mathcal{P}=p\sim\psi_{p},$
(intercept) $\alpha\sim\Gamma,$
(error variance) $\sigma^{2}\sim\Delta,$
(observed data) $(X,Y)_{1:n}\mid b,\tau,\alpha,\sigma^{2}\sim P_{b,\tau,\alpha,\sigma^{2}}(X,Y)\quad\text{i.i.d.},$
where $\pi$, $\phi_{p}$, $\psi_{p}$, $\Gamma$ and $\Delta$ are probability measures on $\mathbb{N}$, $\mathbb{R}^{p}$, $0,1]^{p}$, $\mathbb{R}$ and $\mathbb{R}^{+}_{0}$, respectively. Lastly, define the random variable $\theta=(b,\tau,\alpha,\sigma^{2})$, which takes values in $\Theta$, and denote by $\Pi$ the prior distribution on $\theta$ implied by the model in (3.2).
3.1 Consistency via Doob’s theorem
It turns out that, under very general conditions, the posterior distribution is always consistent at almost every value of $\theta_{0}$ with respect to the measure induced by the prior (Doob, 1949). We state this classical result in the general setting introduced earlier.
@@ 3.1 Consistency via Doob’s theorem
Let the state space $\mathcal{X}$ and the parameter space $\Theta$ be complete separable metric spaces, endowed with their respective Borel sigma-algebras. If $\theta\mapsto P_{\theta}$ is one-to-one and $\theta\mapsto P_{\theta}(A)$ is measurable for all measurable sets $A\subseteq\mathcal{X}$, then the posterior distribution is consistent at $\Pi$-almost all values of $\Theta$. That is, there exists $\Theta_{*}\subseteq\Theta$ such that $\Pi(\Theta_{*})=1$ and for all $\theta_{0}\in\Theta_{*}$, if $X_{1},X_{2},\ldots\sim P_{\theta_{0}}$ i.i.d., then for any neighborhood $B$ of $\theta_{0}$ we have
$\lim_{n\to\infty}\Pi_{n}(\theta\in B\mid X_{1},\dots,X_{n})=1\quad P_{0}^{\infty}-\text{a.s.}$
@@
At this point we can follow an approach very similar to the one in Miller (2023), where posterior consistency is established through Doob’s theorem in a mixture model with an unknown number of components, which has an infinite-dimensional parameter space that factorizes in the same way as (3.1). Considering that we will only be interested in small balls around the true value of the parameter, we can define a bounded metric for $\theta,\theta^{\prime}\in\Theta$ by
@@
$d_{\Theta}(\theta,\theta^{\prime})=\begin{cases}\min\left\{\|\theta-\theta^{\prime}\|,1\right\},\quad&\text{if }p(\theta)=p(\theta^{\prime}),\\ 1,\quad&\text{otherwise}.\end{cases}$ (3.3)
@@
Since each $\Theta_{p}$ is itself a complete separable metric space with the inherited Euclidean norm, Proposition A.1 in Miller (2023) ensures that $(\Theta,d_{\Theta})$ is a complete separable metric space. Note that $P_{\theta}(X,Y)$ is invariant under permutations of the component labels $b$ and $\tau$, so in an identifiable model we can only show consistency in the parameter space up to one such permutation. To that effect, let $S_{p}$ denote the set of permutations of $\{1,\dots,p\}$, and for $\theta\in\Theta_{p}$ and any $\nu\in S_{p}$, denote by $\theta[\nu]$ the result of applying the permutation $\nu$ to the component labels of $\theta$. That is, if $\theta=(\beta_{1},\dots,\beta_{p},t_{1},\dots,t_{p},\alpha,\sigma^{2})$, then $\theta[\nu]=(\beta_{\nu_{1}},\dots,\beta_{\nu_{p}},t_{\nu_{1}},\dots,t_{\nu_{p}},\alpha,\sigma^{2})$. Next, for $\theta_{0}\in\Theta_{p}$ and $\varepsilon>0$ define the neighborhood $\tilde{B}(\theta_{0},\varepsilon)=\bigcup_{\nu\in S_{p}}\{\theta\in\Theta:d_{\Theta}(\theta_{0}[\nu],\theta)<\varepsilon\}$, which consists of all parameters that are within $\varepsilon$ of some permutation of (the component labels of) $\theta_{0}$. Now, for identifiability to hold, we need to place some restrictions on the prior.
Condition 4 (Identifiability constraints).
@@
Both assumptions can be interpreted as a way of pursuing parsimony in the model, aiming for as few components as possible. In practical and computational terms, we can think of $\delta$ as the machine precision number, so that virtually all continuous prior distributions satisfy the associated condition when implemented numerically in a computer. With this setup in mind, we are ready to state our own consistency result.
@@
Suppose that Condition 4 holds and the covariance function $K$ of the underlying process $X$ is strictly positive definite. Then, there exists $\Theta_{*}\subseteq\Theta$ such that $\Pi(\theta\in\Theta_{*})=1$ and for all $\theta_{0}\in\Theta_{*}$, if $(X,Y)_{1:\infty}\sim P_{\theta_{0}}(X,Y)$ i.i.d., then for all $\varepsilon>0$
$\lim_{n\to\infty}\Pi_{n}(\theta\in\tilde{B}(\theta_{0},\varepsilon)\mid(X,Y)_{1:n})=1\quad P_{0}^{\infty}(X,Y)-\text{a.s.}$
@@
$\lim_{n\to\infty}\Pi_{n}(\mathcal{P}=p(\theta_{0})\mid(X,Y)_{1:n})=1\quad P_{0}^{\infty}(X,Y)-\text{a.s.}$